![[ ESO ]](/images/eso-logo.gif) |
 |
La Silla - Science Operation Department

History of changes at the 2.2-m telescope
|
|
Detailed schedule of changes
The schedule of changes below details the dates at which modifications were
carried out at WFI which can affect the flat fields and other properties
of the instrument. Data in one configuration should not be reduced with
calibrations from other. This is particularly true for flat fields and for
zero point maps.
- Configuration 12: (26 Dec 2005) WFI was warmed-up and cooled down. Decontaminated CCDs.
- Configuration 11: (14 September 2005) WFI CCDs were contaminated. New configuration after detox in this date.
- Configuration 10: (23 June 2005) realuminization of M1
- Configuration 9: (16 March 2005 to present date) baffle 5 plus FEROS ADC in FEROS adapter
- Configuration 8: (5 March 2005 -- 15 March 2005) baffle 3 without FEROS adapter.
- Configuration 7: (18 April 2004-5 March 2005) baffle4 plus added velvet to eliminate
unwanted reflection spots: baffle5. Covered selected annular
segments of the walls surrounding the light path
with velvet.
- Configuration 6: (XXX date) baffle3 plus FEROS, plus extra
baffle ring in FEROS adapter: baffle4
- Configuration 5: (XXX date) baffle3 plus FEROS adapter
- Configuration 4: (Beginning of 2003) Improved baffle: baffle3, no vignetting.
- Configuration 3: (12 December 2002) Rings taken out (but spacer in),
return to original baffle1.
- Configuration 2: (18 September 2002) Before FEROS, rings
added to baffle: baffle2. The rings were found to cause
a slight vignette but were on the telescope for a
couple of months. At the same time a spacer was
added between WFI and the telescope (20 mm).
Now the distance of M1 and M2 is not the same
as before. This should add negligible aberrations
but might affect precision astrometry.
- Configuration 1: Beginning of WFI until 18 September 2002) Original baffle: baffle1
By far the large effect occurred at Configuration 7, as the
velvet appears to have reduced the light concentration
problem by a factor of two. There are other changes not noted
above as they can be found in the headers of the images: the
CCD operation temperature was has varied between
167K and 169K in several occasions.